# A Novel Approach to Balance Convenience and Nutrition in Meals With Long-Term Group Recommendations and Reasoning on Multimodal Recipes and its Implementation in BEACON

### v1

@@ Appendix B Boosted Bandit Algorithm for Meal Recommendation
Here we describe how we utilized contextual bandits and reinforcement learning to recommend meals to users. In our application of the boosted-bandit algorithm (Kakadiya, Natarajan, and Ravindran 2021), contextual bandits automatically extract and learn user preferences and provide highly personalized meal recommendations. Given a set of users with their dietary preferences($U$) as well as the set of all recipes in R3 format($R$), we create predicate logic pairs of the form preference(user_5, negative_nuts) and item(food_18, has_nuts), indicating that the $user\_5$ has a negative preference for items with nuts and that food item 18 in $R$ has nuts. Following this, we also create negative and positive predicate logic pairs of the form recommendation(user_18, food_22), to signify if a recommendation is a positive or negative recommendation to the user based on their aligning preferences. We then split these predicate logic pairs into train and test sets and train the boosted-bandits  (Kakadiya, Natarajan, and Ravindran 2021). Then, we test the bandits on the test set, and for each recommendation(user, item) pair, we receive a corresponding probability, which represents how likely the user is to prefer that item. We then construct a recommendation, $\mathcal{MP}$, by utilizing the users’ preferred recommendation output (see Problem Formulation for details) and the highest probability items.

### v2

@@ Appendix B Boosted Bandit Algorithm for Meal Recommendation
Here we describe how we utilized contextual bandits and reinforcement learning to recommend meals to users. In our application of the boosted-bandit algorithm (Kakadiya, Natarajan, and Ravindran 2021), contextual bandits automatically extract and learn user preferences and provide highly personalized meal recommendations. Given a set of users with their dietary preferences($U$) as well as the set of all recipes in R3 format($R$), we create predicate logic pairs of the form preference(user_5, negative_nuts) and item(food_18, has_nuts), indicating that the user_5user\_5 has a negative preference for items with nuts and that food item 18 in $R$ has nuts. Following this, we also create negative and positive predicate logic pairs of the form recommendation(user_18, food_22), to signify if a recommendation is a positive or negative recommendation to the user based on their aligning preferences. We then split these predicate logic pairs into train and test sets and train the boosted-bandits  (Kakadiya, Natarajan, and Ravindran 2021). Then, we test the bandits on the test set, and for each recommendation(user, item) pair, we receive a corresponding probability, which represents how likely the user is to prefer that item. We then construct a recommendation, $\mathcal{MP}$, by utilizing the users’ preferred recommendation output (see Problem Formulation for details) and the highest probability items.
@@ Appendix C BEACON Implementation
System Figures
Figures 11, 12, and 13 show the meal, food, and ingredient views of BEACON’s interactive meal calendar, respectively.
Figure 11: Calendar-based meal visualization screen on the Meal View.
Figure 12: Calendar-based meal visualization screen on the Food Item View.
Figure 13: Calendar-based meal visualization screen on the Ingredient View.